# User Location Disclosure Fails to Deter Overseas Criticism but Amplifies Regional Divisions on Chinese Social Media

### v2

@@ 5 Conclusion
To illustrate, we reanalyze the main result from Figure 2(a), which shows a post-policy increase in comment volume from overseas users on international topics. Figure S12 replicates this analysis using the five alternative bandwidths. With a narrower bandwidth (8 hours; Figure 12(a)), the smoothing line appears more variable, but the pattern remains: comments on international posts rise sharply, while those on non-international posts remain stable. This finding holds consistently with wider bandwidths of 24 hours (Figure 12(b)), 72 hours (Figure 12(c)), 96 hours (Figure 12(d)), and 120 hours (Figure 12(e)).
@@
Figure S13 replicates Figure 2(b) using alternative bandwidths. Across all specifications, comments from domestic users consistently decline only on non-international posts, mirroring the pattern observed with the cross-validated bandwidth.
@@
Figure S14 tests the robustness of our main finding in Figure 2(c) by examining comment dynamics on non-international posts under varying bandwidths (8, 24, 72, 96, and 120 hours). Across all bandwidth choices, the decline in comments on local issues after the user location disclosure policy remains consistent, confirming that the observed pattern is not sensitive to bandwidth selection and reinforcing the reliability of our results.
@@
Figure S15 examines the decline in comments on local topics by commenter origin, replicating Figure 2(e) with varying bandwidths. Across all bandwidth choices, the estimated reduction in out-of-province comments remains robust, confirming that the user location disclosure policy disproportionately affected non-local engagement. This reinforces our conclusion that the policy significantly curtailed participation from geographically distant users.
@@
Figure S16 extends the analysis by specifically focusing on critical comments and their geographical origin (Figure 2(f)), further validating the robustness of the findings. The consistent decline observed in out-of-province critical comments highlights that the user location disclosure significantly affects negative or dissenting voices from geographically distant commenters, reinforcing the broader implications of the policy on public discourse.
@@
Figure S17 replicates the analysis in Figure 3(a) using multiple bandwidths to assess the robustness of the observed rise in regional discrimination following the user location disclosure. Across all bandwidth choices, the increase in discriminatory replies remains robust, supporting our finding that the policy exacerbated regional antagonism in online interactions.
@@
Table S8: Treatment Effect Results Table (Time Placebo)
@@
Figure 2(a) Topic: International #Comments 0.369 -0.226 0.179
@@
Figure 2(a) Topic: Non-International #Comments -0.011 0.034 -0.036
@@
Figure 2(b) Topic: International #Comments 0.407 -0.487 -0.186
@@
Figure 2(b) Topic: Non-International #Comments -4.110 -0.504 -1.149
@@
Figure 2(c) Topic: Non-Local #Comments -1.399 -0.154 -0.627
@@
Figure 2(c) Topic: Local #Comments -7.794 1.872 -0.288
@@
Figure 2(e) User: In-Province #Comments -1.710 -0.406 0.137
@@
Figure 2(e) User: Out-of-Province #Comments -5.665 0.761 0.324
@@
Figure 2(e) User: Oversea #Comments -0.073 -0.070 -0.032
@@
Figure 2(f) User: In-Province #Critical Comments -0.572 0.070 -0.178
@@
Figure 2(f) User: Out-of-Province #Critical Comments -3.591 1.902 -2.622
@@
Figure 2(f) User: Oversea #Critical Comments -0.030 -0.042 -0.057
@@
In Figure 3(a), we documented an increase in regionally discriminatory replies after the implementation of the user location disclosure policy. To unpack this trend, Figure S26 disaggregates discriminatory replies by interaction type. Panel 26(a) shows that the rise is largely driven by inter-group interactions, while within-group discrimination remains stable. Panel 26(b) further reveals that most inter-group discrimination comes from local users targeting non-local users, underscoring the role of geographic identity in shaping online hostility.

### v3

@@ 4.2 Causal Identification
To illustrate, we reanalyze the main result from Figure 1a, which shows a post-policy increase in comment volume from overseas users on international topics. Figure S12 replicates this analysis using the five alternative bandwidths. With a narrower bandwidth (8 hours; Figure 12(a)), the smoothing line appears more variable, but the pattern remains: comments on international posts rise sharply, while those on non-international posts remain stable. This finding holds consistently with wider bandwidths of 24 hours (Figure 12(b)), 72 hours (Figure 12(c)), 96 hours (Figure 12(d)), and 120 hours (Figure 12(e)).
@@
Figure S13 replicates Figure 1b using alternative bandwidths. Across all specifications, comments from domestic users consistently decline only on non-international posts, mirroring the pattern observed with the cross-validated bandwidth.
@@
Figure S14 tests the robustness of our main finding in Figure 1c by examining comment dynamics on non-international posts under varying bandwidths (8, 24, 72, 96, and 120 hours). Across all bandwidth choices, the decline in comments on local issues after the user location disclosure policy remains consistent, confirming that the observed pattern is not sensitive to bandwidth selection and reinforcing the reliability of our results.
@@
Figure S15 examines the decline in comments on local topics by commenter origin, replicating Figure 1e with varying bandwidths. Across all bandwidth choices, the estimated reduction in out-of-province comments remains robust, confirming that the user location disclosure policy disproportionately affected non-local engagement. This reinforces our conclusion that the policy significantly curtailed participation from geographically distant users.
@@
Figure S16 extends the analysis by specifically focusing on critical comments and their geographical origin (Figure 1f), further validating the robustness of the findings. The consistent decline observed in out-of-province critical comments highlights that the user location disclosure significantly affects negative or dissenting voices from geographically distant commenters, reinforcing the broader implications of the policy on public discourse.
@@
Figure S17 replicates the analysis in Figure 2a using multiple bandwidths to assess the robustness of the observed rise in regional discrimination following the user location disclosure. Across all bandwidth choices, the increase in discriminatory replies remains robust, supporting our finding that the policy exacerbated regional antagonism in online interactions.
@@
Table S8: Placebo Estimates
@@
Figure 1a Topic: International #Comments 0.369 -0.226 0.179
@@
Figure 1a Topic: Non-International #Comments -0.011 0.034 -0.036
@@
Figure 1b Topic: International #Comments 0.407 -0.487 -0.186
@@
Figure 1b Topic: Non-International #Comments -4.110 -0.504 -1.149
@@
Figure 1c Topic: Non-Local #Comments -1.399 -0.154 -0.627
@@
Figure 1c Topic: Local #Comments -7.794 1.872 -0.288
@@
Figure 1e User: In-Province #Comments -1.710 -0.406 0.137
@@
Figure 1e User: Out-of-Province #Comments -5.665 0.761 0.324
@@
Figure 1e User: Oversea #Comments -0.073 -0.070 -0.032
@@
Figure 1f User: In-Province #Critical Comments -0.572 0.070 -0.178
@@
Figure 1f User: Out-of-Province #Critical Comments -3.591 1.902 -2.622
@@
Figure 1f User: Oversea #Critical Comments -0.030 -0.042 -0.057
@@
In Figure 2a, we documented an increase in regionally discriminatory replies after the implementation of the user location disclosure policy. To unpack this trend, Figure S26 disaggregates discriminatory replies by interaction type. Panel 26(a) shows that the rise is largely driven by inter-group interactions, while within-group discrimination remains stable. Panel 26(b) further reveals that most inter-group discrimination comes from local users targeting non-local users, underscoring the role of geographic identity in shaping online hostility.